Future Value of Annuity Due


1. Jade Savings and Loan is paying 6% interest compounded quarterly. Find the future value of P1,000, deposited at the beginning of every 3 months, for 5 years.

2. Bela invested P2,500 at the beginning of every 3-month period, for 5 years, at 8% interest compounded quarterly. How much is Bela investment worth after 5 years?

3. If Depot Hardware Store invested P1,000 at the beginning of each month at 12% compounded monthly, how much would be in the account after 5 years?

4. Find the future value of an annuity due of P7,000 payable at the beginning of each quarter for six years if the interest rate is 14% compounded quarterly.

5. If P6,000 is deposited at the beginning of each month in a bank that pays 12% interest compounded monthly, what is the final value at the end of three years and four months?


1
Expert's answer
2021-05-07T09:08:55-0400

and t is time.

"1.) A=1000\times (1-(1.015)^{-20})/.015=17,427.55"


"2.) A=2,500\times (1-(1.02)^{-20})/.02=41,694"


3.) "2.) A=1000\times (1-(1.01)^{-60})/.01=45,405"


4.) "A=7000\times (1-(1.035)^{-24})/.035=116,343"


5.) "A=6000\times (1-(1.01)^{-40})/.01=198,978"